OpenStudy (anonymous):

Determine whether the equation is an identity or not an identity. (cos theta) / (sin theta)(tan theta) = 1

OpenStudy (whpalmer4):

Is that supposed to be \[\frac{\cos \theta}{\sin \theta} * \tan\theta\]?

OpenStudy (anonymous):

no its just a true or false question and its exactly like that.

OpenStudy (anonymous):

cot theta * tan theta is 1 yes

OpenStudy (loser66):

hello futuremathprofessor, please help